# Hierarchical Latent Class Models for Mortality Surveillance Using Partially Verified Verbal Autopsies

### v1

@@ 2.4 Stratification and structured priors
A main challenge when considering fine stratification of a population is that inevitably some sub-populations will contain only a small number of deaths. Previous work involving cause-of-death assignment in multiple populations all assume independent priors for $\pi^{(g)}$ (see e.g., McCormick et al.,, 2016; Li et al.,, 2024; Wu et al.,, 2024). For mortality monitoring of small sub-populations, however, it is usually more appropriate to borrow information across related sub-populations in order to improve the stability and interpretability of estimated prevalence. Similar ideas have been extensively studied for survey data in the context of small area estimation (e.g., Rao and Molina,, 2015; Gao et al.,, 2021), but it has not been explored in VA analysis. Unlike the traditional small area estimation literature, the main target of interest in VA analysis, the cause of death, cannot be directly collected and needs to be estimated instead. Nevertheless, we adopt similar ideas to leverage the structured relationship of sub-populations and directly borrow information across strata.
To fix notation, here we consider sub-populations defined by three stratification variables. Let $S_{i}\in\{1,2\}$ indicate sex (1 = female and 2 = male), $T_{i}\in\{1,2,...,T\}$ indicate the time period and $A_{i}=\{1,2,...,A\}$ indicate the age group of the $i$-th death respectively. We reparameterize the nested latent class model with
$p(Y_{i}=1)$ $=\pi_{s[i],t[i],a[i]},$ (4)
$p(Z_{i}=k\mid Y_{i}=c)$ $=\lambda_{ck}^{(s[i],t[i],a[i])}.$ (5)
@@
$\pi_{sta}=\mbox{logit}^{-1}(\mu+\alpha^{\texttt{male}}\mathbbm{1}_{s=1}+\alpha^{\texttt{time}}_{t}+\alpha^{\texttt{age}}_{a}+\epsilon_{sta}).$ (6)
We model $\mu$ and $\alpha^{\texttt{male}}$ as fixed effects with independent $N(0,100)$ priors. For the time and age effects, We consider the following three prior specifications:
@@
The three prior specifications differ in the amount of information shared across strata. In the fixed effect model, no information is shared except for the additive structure of the time and age effect. The independent random effect model further shrinks these effects to the common mean. The random walk model shrinks these effects so strata with similar ages or time periods will have similar prevalence.
To assess the benefits of incorporating structured priors, we also compare the three models above with an unstructured baseline model, where $\pi_{sta}\overset{iid}{\sim}\mbox{Beta}(1,1)$. This serves as a benchmark for evaluating the impact of the structured priors on the model’s performance. The unstructured prior is the standard practice in the literature and practice of VA analysis (McCormick et al.,, 2016; Li et al.,, 2023). Lastly, when the target of inference is only the overall prevalence, one may consider an unstratified baseline model, where $\pi_{sta}=\pi_{0}\sim\mbox{Beta}(1,1)$ for all strata. As we have discussed in Section 2.1, this unstratified model can lead to biased prevalence estimates if the verification mechanism changes over variables not accounted for by the model (e.g., time period). We will demonstrate such bias in simulation and real data analysis in Section 4.2.
@@ 3 Posterior inference
    $p(Y_{i}=c\mid\bm{X_{i}},\bm{\pi},\bm{\phi},\bm{\lambda})\propto\pi_{s[i],t[i],a[i]}^{c}(1-\pi_{s[i],t[i],a[i]})^{1-c}\sum_{k=1}^{K}\lambda_{ck}^{(s[i],t[i],a[i])}\prod_{j=1}^{p}\phi_{ckj}^{X_{ij}}(1-\phi_{ckj})^{1-X_{ij}}.$
@@
    where $\bm{M}=\bm{V}\bm{P}^{T}(\bm{z}-\bm{n}/2)$ and $\bm{V}=(\bm{\Sigma}^{-1}+\bm{P}^{T}diag(\bm{\omega})\bm{P})^{-1}$.
@@ 4.1 Simulated data
We consider two types of verification mechanisms in the simulation study. In case (i), the verification process depends on observed symptoms and covariates $X$ and $D$. In case (ii), the verification process also depends on the unobserved cause of death $Y$. In the first case, unbiased estimation of prevalence is possible as Assumption 1 is satisfied. In the second case, the verification process is not conditionally ignorable and can lead to bias in parameter estimation. This situation can arise in practice when the verification process depends on variables associated with the cause of death but is unavailable to the data analysts. More specifically,
$p(L_{i}\mid X_{i}=\bm{x},A_{i}=a,T_{i}=t,Y_{i}=y)=\mbox{logit}^{-1}(a^{\texttt{time}}_{t}+a^{\texttt{age}}_{a}+\bm{b}_{ta}^{T}\bm{x}+c_{1}y+c_{2}(1-y))$
where $c_{1}=c_{2}=0$ in case (i), and $c_{1}\sim\mbox{Unif}(-0.4,0)$, $c_{2}=-c_{1}$ in case (ii). For the choice of the coefficients in the verification mechanism, we consider the more realistic situations where we under-sample time periods with more deaths and over-sample time periods with fewer deaths. We also over-sample the first two and last two age groups to reflect situations where the mortality patterns of younger and elder groups are of greater interest. For the other symptoms, we allow them to be weakly associated with the selection probability in a time-varying fashion. Specifically, we let
@@
To mimic the true prevalence through time and age groups, we generate the true prevalence $\pi_{sta}$ with the additive polynomial trend in both time and age dimensions. The simulated prevalence is shown as the black dots in Figure 5. The cause-of-death $Y$ and symptoms $X$ are generated according to the latent class model described in Section 2 with $K=10$. We consider $q=10$, $T=10$, and $A=8$. The sample size for each sub-population is balanced with $n_{sta}=100$. We generate $50$ synthetic datasets and fit all models with $K=10$.
We first illustrate the bias induced by not accounting for the verification process discussed in Section 2.1. Here we compare the four models stratified by sex, age, and time, with the unstratified baseline model described in Section 2.4. This baseline corresponds to the naive adoption of VA cause-of-death assignment algorithms without considering the time-varying verification probability. It should be noted that the unstratified baseline model also includes the latent class model component and is more flexible than the VA models currently used in practice that assume conditional independence of symptoms McCormick et al., (2016); Byass et al., (2019). Since the unstratified model aims to estimate the overall prevalence in the population, we aggregate the estimated stratum-specific prevalence in the other four models to the population-level prevalence by
@@
We take the posterior mean as our point estimates $\hat{\bm{\pi}}$ and evaluate the bias compared to the true prevalence, $\hat{\bm{\pi}}-\bm{\pi}$. Panel (a) of Figure 4 shows that the unstratified model leads to the largest bias across the simulated datasets in case (i). The three models with structured prior perform similarly and achieve much smaller bias compared to the unstructured model. Panel (c) of Figure 4 shows similar patterns for case (ii). It shows that the models that partially account for the verification process are also more robust when the conditional ignorability assumption is violated.
Next, we consider the bias from ignoring the heterogeneity of symptom distributions across strata, discussed in Sec 2.3. We consider the comparison between the models stratified by sex, age, and time, and simplified models stratified by time only. In the latter case, we include sex and age group as binary dummy variables in $X$, and remove the sex- and age-effect from the model for $\bm{\pi}$. For the fully stratified model, we obtain the time-varying overall prevalence using a similar aggregation step as before,
@@
Panel (b) and (d) of Figure 4 compare the bias of $\hat{\pi}^{(t)}$ based on the model with partial and full stratification, under the unstructured baseline and the random walk model. The fixed effect and models also show similar patterns. It can be observed that the fully stratified models are able to achieve consistently lower bias in both case (i) and (ii), due to its flexibility to incorporate stratum-specific symptom distribution.
Figure 4: Bias of posterior mean prevalence estimates over $50$ simulated datasets. The top row corresponds to simulation case (i) and the bottom row corresponds to simulation case (ii). Panel (A) and (C) evaluate the bias for population-level overall prevalence estimates and panel (B) and (D) evaluate the bias for time-varying overall prevalence estimates.
Finally, we evaluate the effect of the structured prior on the stratum-specific prevalence estimates. Figure 5 illustrates the posterior means and 95% credible intervals for the prevalence estimates derived from a specific synthetic dataset in case (i). The unstructured baseline model yields prevalence estimates that exhibit notably higher variability over time and larger uncertainty. In contrast, the random walk model yields smoother estimates with narrower credible intervals, and generally captures the true value of prevalence more accurately.
Figure 5: Posterior mean and 95% credible intervals of the estimated prevalence in one simulated dataset for different age groups, sex, and months. The true prevalence is indicated by the black dots.
4.2 Brazil COVID-19 surveillance data
We evaluate our methods on a flu syndrome surveillance dataset collected in Brazil from January to October, 2021. All deaths in this database had severe acute respiratory syndrome by COVID, other respiratory diseases or chronic diseases. The dataset contains the final cause of death for $411,491$ deaths. We consider $p=14$ binary indicators for each death, including symptoms such as fever, vomiting, loss of taste and smell, etc. In our analysis, we exclude the information on vaccine, PCR and antigen test results as they are largely missing. Within this population of deaths suspected of COVID-19, the proportion of deaths with the final classification of being COVID-19 related increases from January to April and declines afterward. We consider eight age groups, where the first includes deaths between 0 and 30 years old and the subsequent groups are defined by ten-year increments. We let the first age group span 30 years because deaths under $20$ years old are rare in this population. Figure 6 shows the true prevalence, i.e., the fraction of deaths due to COVID-19 among these suspected deaths, under each sub-population. The sex-specific trends differ from each other only mildly. Figure 7 illustrates that the sample sizes across sub-populations are notably imbalanced, with particularly small samples observed in January and October, as well as among the younger and elderly age groups.
Figure 6: Fraction of deaths due to COVID-19 among the deaths suspected of COVID-19 across different age and sex strata over time.
We randomly sample $50\%$ observations within each sex $s$, month $t$ and age group $a$, while keeping the proportion of COVID-19 related death the same as the prevalence in the full population. We assume a verification mechanism that depends on time, age, and a random subset of symptoms, with the same coefficient as described in the simulation study,
@@ 4.2 Brazil COVID-19 surveillance data
We generate $50$ synthetic datasets and fit all models with $K=10$. We run the MCMC for $8000$ iterations with $3000$ iterations as burn-in.
Figure 8 shows the posterior mean and 95% credible interval of the stratum-specific prevalence in one of the synthetic datasets, estimated by the four stratified models. Stronger effect of smoothing can be observed in the two random effect models, especially in the last two time periods when the total number of deaths is smaller.
Figure 8: Posterior mean and 95% credible intervals of the estimated prevalence using the four models for different age group, sex, and months, based on one synthetic dataset resampled from the COVID-19 surveillance data from Brazil. The true prevalence is indicated by the black dots.
@@
where $X\sim F$ and $X^{\prime}\sim F$ are two independent random variables, and $x$ is the true prevalence. CRPS is positive, where values closer to 0 indicate better prediction. We compute the CRPS difference between the baseline and each of the three structured models. Positive differences indicate that the structured models outperform the baseline. Figure 9 shows the CRPS differences for the three models, compared to the unstructured baseline model, across all sub-populations in the $50$ datasets. As expected, we can observe improved CRPS from all three structured models when the sample size is small and the proportion of unverified deaths is high.
@@
Figure 9: Improvement in CRPS compared to the unstructured baseline model for the three structured models, in terms of difference in CRPS. Each dot represents one sub-population in one simulated dataset, arranged by the sample size of the sub-population, and the proportion of unverified labels, i.e., the proportion of deaths without a cause, in the sub-population. The red line is the smoothed conditional mean for the points.
@@
Figure 10 shows the similar pattern of bias as illustrated in the simulation study. The unstructured baseline model results in the largest bias for the population-level overall prevalence, due to not accounting for the time-varying verification mechanism. Figure 11 shows that for all models, accounting for the heterogeneity of symptom distributions over sex and age also greatly improves prevalence estimation.
@@
Figure 10: Distribution of bias of overall prevalence estimation over $50$ synthetic datasets resampled from the COVID-19 surveillance data from Brazil, under the population-level model and the four stratified models.
@@
Figure 11: Bias of time-varying prevalence estimation over $50$ synthetic datasets resampled from the COVID-19 surveillance data from Brazil, under the four models stratified by sex, age, and time (top row) and the models stratified by time only (bottom row).
@@
Finally, the proposed latent class model framework also allows us to gain more insight into how the data distribution shifts over time by examining the estimated latent parameters. Figures 12 visualizes the heatmap of latent symptom profiles and the corresponding weights across different strata under the random walk model in one synthetic data. The latent classes are ordered by the expected number of symptoms under the symptom profiles of each class. We can observe distinct clusters of symptoms, such as fever and cough; dyspnea, respiratory discomfort, and low $O_{2}$ saturation; loss of taste and loss of smell, that are more likely to occur in the same latent symptom profiles, which seem to correspond to symptoms with shared mechanism. The estimated weights $\bm{\lambda}$ illustrate the changes in abundance of these latent classes over age, sex, and time. For example, among the deaths related to COVID-19, the weights of the fifth latent class are much larger in male deaths under 60 years old, compared to other demographic groups, whereas the sixth latent class has larger weights among deaths above 60 years old. Both symptom profiles include high probabilities of dyspnea, respiratory discomfort, and low $O_{2}$ saturation. The main difference in these two latent symptom profiles is that the former includes higher probability of fever and cough, whereas the latter includes low probability of fever and cough, but higher probability of having at least one risk factors. These patterns are consistent with the explorative analysis in Figure 2.
@@
Figure 12: Posterior mean of latent parameters estimated in one synthetic dataset resampled from the COVID-19 surveillance data from Brazil. Panel (a): symptom profiles $\bm{\phi}$ by latent class for deaths related to COVID-19, ordered by the expected number of symptoms. Panel (b): latent class weights $\bm{\lambda}$ over sex, age, and time for deaths related to COVID-19. Panel (c) and (d) are the same parameters for deaths not related to COVID-19.

### v3

@@ 3.3 Stratification and structured priors
A main challenge when considering fine stratification of a population is that inevitably some sub-populations will contain only a small number of deaths. Previous work involving cause-of-death assignment in multiple populations all assume independent priors for $\pi^{(g)}$ (see e.g., McCormick et al.,, 2016; Li et al.,, 2024; Wu et al.,, 2024). For mortality monitoring of small sub-populations, however, it is usually more appropriate to borrow information across related sub-populations in order to improve the stability and interpretability of estimated CSMFs. Similar ideas have been extensively studied for survey data in the context of small area estimation (e.g., Rao and Molina,, 2015; Gao et al.,, 2021). Small area estimation methods have not been explored in VA analysis. Unlike the traditional small area estimation literature, the outcome of interest, i.e., the cause of death, cannot be directly collected and needs to be estimated instead. However, the key idea of leveraging the structured relationship of sub-populations and borrowing information across strata applies naturally to mortality surveillance using VAs.
To fix notation, here we describe the model where sub-populations $D$ are defined by three stratification variables. Let $S_{i}\in\{1,2\}$ indicate sex (1 = female and 2 = male), $T_{i}\in\{1,2,...,T\}$ indicate the time period, and $A_{i}=\{1,2,...,A\}$ indicate the age group of the $i$-th death respectively. The hierarchical latent class model with $D_{i}=(S_{i},T_{i},A_{i})$ can be written as
$p(Y_{i}=1\mid D_{i})$ $=\pi_{s[i],t[i],a[i]},$ (4)
$p(Z_{i}=k\mid Y_{i}=c,D_{i})$ $=\lambda_{ck}^{(s[i],t[i],a[i])},\;\;\;\;c=0,1.$ (5)
@@
$\pi_{sta}=\mbox{logit}^{-1}(\mu+\alpha^{\texttt{male}}\mathbbm{1}_{s=1}+\alpha^{\texttt{time}}_{t}+\alpha^{\texttt{age}}_{a}+\epsilon_{sta}),$ (6)
where $\mathbbm{1}_{s=1}$ denotes the indicator function that takes value of $1$ when $s=1$ and 00 otherwise. We model $\mu$ and $\alpha^{\texttt{male}}$ as fixed effects with independent $N(0,100)$ priors. We let $\epsilon_{sta}\overset{iid}{\sim}N(0,\sigma_{\epsilon}^{2})$ be an unstructured interaction term that captures the deviation from the main additive decomposition. For the time and age effects, We consider the following three prior specifications:
@@
The three prior specifications differ in the amount of information shared across strata. In the fixed effect model, no information is shared except for the additive structure of the time and age effect. The independent random effect model further shrinks these effects to the common mean. The random walk model shrinks the first order difference in these effects so that strata with similar ages or time periods will have similar CSMFs.
@@ 4 Posterior inference
    $p(Y_{i}=c\mid\bm{X_{i}},\bm{\pi},\bm{\phi},\bm{\lambda})\;\propto\;$ $\pi_{s[i],t[i],a[i]}^{c}(1-\pi_{s[i],t[i],a[i]})^{1-c}$
    ×$\times\sum_{k=1}^{K}\lambda_{ck}^{(s[i],t[i],a[i])}\prod_{j=1}^{p}\phi_{ckj}^{X_{ij}}(1-\phi_{ckj})^{1-X_{ij}}.$
@@
    where $\bm{M}=\bm{V}\bm{P}^{T}(\bm{z}-\bm{n}/2)$ and $\bm{V}=(\bm{\Sigma}^{-1}+\bm{P}^{T}\mbox{diag}(\bm{\omega})\bm{P})^{-1}$.
@@ 5.1 Simulated data
We start with fully synthetic data simulation. To mimic the true CSMFs across time and age groups, we generate the true $\pi_{sta}$ with the additive polynomial trend in both time and age dimensions. The simulated CSMFs are shown as the black dots in Figure 7. The cause of death $Y$ and symptoms $X$ are generated according to the latent class model described in Section 3 with $K=10$. In the simulation, we let $p=10$, $T=10$, and $A=8$. The sample size for each sub-population is balanced with $n_{sta}=100$. We generate $50$ synthetic datasets and fit all models with $K=10$.
We consider the following verification mechanism, where the stratification variables include age of the deceased and the time period of death, i.e., $D_{i}=(A_{i},T_{i})$. We let
$p(L_{i}\mid X_{i}=\bm{x},A_{i}=a,T_{i}=t,Y_{i}=y)=\mbox{logit}^{-1}(a^{\texttt{time}}_{t}+a^{\texttt{age}}_{a}+\bm{b}_{ta}^{T}\bm{x}+c_{1}y+c_{2}(1-y)).$
We consider two types of verification mechanisms in the simulation study.
  • Simulation case (i): The verification process depends on observed symptoms $X$ and covariates $D$. We let $c_{1}=c_{2}=0$. Assumption 1 is satisfied in this case.
  • Simulation case (ii): The verification process depends on observed symptoms $X$, covariates $D$, and also the unobserved cause of death $Y$. We let $c_{1}\sim\mbox{Unif}(-0.4,0)$, $c_{2}=-c_{1}$. In this case, the verification process is not conditionally ignorable, and CSMF estimation can be biased. This situation can arise in practice when the verification process depends on variables associated with the cause of death but is unavailable to the data analysts.
For the choice of the coefficients in the verification mechanism, we consider the more realistic situations where we under-sample time periods with more deaths and over-sample time periods with fewer deaths. We also over-sample the first two and last two age groups to reflect situations where the mortality patterns of the younger and elderly groups are of greater interest. For the other symptoms, we allow them to be weakly associated with the selection probability in a time-varying fashion. Specifically, we let
@@
In addition to the three models discussed in Section 3.3, to assess the benefits of the proposed models, we consider the following alternative models:
  1. 1.
    Time-only model: We include sex and age group as binary dummy variables in $X$, and remove the sex- and age-effect from the model for $\bm{\pi}$. This leads to a simplified model that is stratified by time. This model represents the scenario where the time-varying CSMF is of interest, but the stratification by sex and age is not accounted for in the prior for $\bm{\pi}$. This model ignores the heterogeneity of symptom distributions across different age groups and sexes.
  2. 2.
    Unstructured baseline model: We use independent priors for $\bm{\pi}$ without borrowing information. That is, we let $\pi_{sta}\overset{iid}{\sim}\mbox{Beta}(1,1)$. This model is a simpler alternative to the structured priors proposed in Section 3.3.
  3. 3.
    Unstratified baseline model: In this model, we simply let $\pi_{sta}=\pi_{0}\sim\mbox{Beta}(1,1)$ for all strata. As we have discussed in Section 3.1, this unstratified model can lead to biased CSMF estimates if the verification mechanism changes over variables not accounted for by the model. This baseline is only relevant when the target of inference is the overall CSMF.
It should be noted that in terms of modeling $\bm{\pi}$, the unstratified baseline model is the default approach in the VA literature and routine analysis when only the overall CSMF is of interest, and the unstructured baseline model is the standard approach when multiple sub-populations are of interest (McCormick et al.,, 2016; Li et al.,, 2023). In our study, both baseline models adopt the latent class model framework in modeling $p(X\mid Y)$. Thus, they are already more flexible than the VA models currently used in practice, which assume conditional independence of symptoms (McCormick et al.,, 2016; Byass et al.,, 2019). Given the extensive evidence that the latent class model framework improves from models making the conditional independent assumption (Li et al.,, 2024; Wu et al.,, 2024), we do not further compare our methods to simpler models.
We first illustrate the bias induced by not accounting for the verification process discussed in Section 3.1. Here we compare the three stratified models with structured priors, the unstructured baseline model, and the unstratified baseline model. Since the unstratified model aims to estimate the overall CSMF in the population, we aggregate the estimated stratum-specific CSMFs in the other four models to the population-level CSMF by
@@
We take the posterior mean as our point estimate and evaluate the bias compared to the true CSMF, i.e., $\hat{\pi}-\pi$. Panel (a) of Figure 6 shows that in case (i), the unstratified model leads to the largest bias across the simulated datasets. The three models with structured prior perform similarly and achieve much smaller bias compared to the unstructured model. Panel (c) of Figure 6 shows similar patterns for case (ii). In this case, the proposed models partially account for the verification process and are more robust than the unstratified model even when the missing at random assumption is violated.
Next, we consider the bias from ignoring the heterogeneity of symptom distributions across strata, discussed in Section 3.2. We consider the comparison between the models stratified by sex, age, and time, and the time-only model. For the fully stratified model, we obtain the time-varying overall CSMF using a similar aggregation step as before,
@@
Panels (b) and (d) of Figure 6 compare the bias of $\hat{\pi}^{(t)}$ based on the model with partial and full stratification, under the unstructured baseline and the random walk model. The fixed effect and models also show similar patterns. It can be observed that the fully stratified models are able to achieve consistently lower bias in both cases, due to their flexibility to incorporate stratum-specific symptom distribution.
Figure 6: Bias of posterior mean CSMF estimates over $50$ simulated datasets. The top row corresponds to simulation case (i), and the bottom row corresponds to simulation case (ii). Panels (A) and (C) evaluate the bias for population-level overall CSMF estimates, and panels (B) and (D) evaluate the bias for time-varying overall CSMF estimates.
Finally, we evaluate the effect of the structured prior on the stratum-specific CSMF estimates. Figure 7 illustrates the posterior means and 95% credible intervals for the CSMF estimates derived from a single synthetic dataset in case (i). The unstructured baseline model yields CSMF estimates that exhibit notably higher variability over time with larger uncertainty. In contrast, the random walk model yields smoother estimates with narrower credible intervals and generally captures the actual CSMF more accurately.
Figure 7: Posterior mean and 95% credible intervals of the estimated CSMF in one simulated dataset for different age groups, sex, and months. The true CSMF is indicated by the black dots.
5.2 Brazil COVID-19 analysis
We now evaluate our methods using the Brazil COVID-19 surveillance dataset described in Section 2. We consider the case where $D_{i}=(S_{i},A_{i},T_{i})$, that is the data are stratified by sex, age, and time periods. To create realistic replications, we generated $100$ semi-synthetic datasets through resampling. In each resampled dataset, we randomly sample $50\%$ observations within each sex, month, and age group, while keeping the proportion of COVID-19 related deaths the same as the CSMFs in the full population. We assume a verification mechanism that depends on time, age, and a random subset of symptoms, with the same coefficient as described in the simulation study,
@@ 5.2 Brazil COVID-19 analysis
We fit all models with $K=10$. We run the MCMC for $8000$ iterations with $3000$ iterations as burn-in. Figure 8 shows the posterior mean and 95% credible interval of the stratum-specific CSMF in one resampled dataset, estimated by the four stratified models. Stronger effects of smoothing can be observed in the two random effect models, especially in the last two time periods when the total number of deaths is smaller.
Figure 8: Posterior mean and 95% credible intervals of the estimated CSMF using the four models for different age groups, sex, and months, based on one synthetic dataset resampled from the COVID-19 surveillance data from Brazil. The true CSMFs are indicated by the black dots. Age group 1 corresponds to 0–30 years old. Age groups 2 to 7 correspond to 10-year intervals from 30–40 to 80–90. Age group 8 corresponds to over 90 years old.
@@
where $X\sim F$ and $X^{\prime}\sim F$ are two independent random variables, and $x$ is the true CSMF. CRPS is positive, where values closer to 0 indicate better prediction. We compute the CRPS difference between the baseline and each of the three structured models. Positive differences indicate that the structured models outperform the baseline. Figure 9 shows the CRPS differences for the three models, compared to the unstructured baseline model, across all sub-populations in the $50$ datasets. As expected, we can observe improved CRPS from all three structured models when the sample size is small and the proportion of unverified deaths is high.
@@
Figure 9: Improvement in CRPS compared to the unstructured baseline model for the three structured models, in terms of difference in CRPS. Each dot represents one sub-population in one simulated dataset, arranged by the sample size of the sub-population and the proportion of unverified labels, i.e., the proportion of deaths without a cause, in the sub-population. The red line is the smoothed conditional mean for the points.
@@
Figure 10 shows the similar pattern of bias as illustrated in the simulation study. The unstructured baseline model results in the largest bias for the population-level overall CSMF, due to not accounting for the time-varying verification mechanism. Figure 11 shows that for all models, accounting for the heterogeneity of symptom distributions over sex and age also greatly improves CSMF estimation.
@@
Figure 10: Distribution of bias of overall CSMF estimation over $100$ synthetic datasets resampled from the COVID-19 surveillance data from Brazil, under the population-level model and the four stratified models.
@@
Figure 11: Bias of time-varying CSMF estimation over $100$ synthetic datasets resampled from the COVID-19 surveillance data from Brazil, under the four models stratified by sex, age, and time (top row) and the models stratified by time only (bottom row).
@@
Finally, the proposed latent class model framework also allows us to gain more insight into how the data distribution shifts over time by examining the estimated latent parameters. Figure 12 visualizes the heatmap of latent symptom profiles and the corresponding weights across different strata under the random walk model in one synthetic dataset. The latent classes are ordered by the expected number of symptoms under the symptom profiles of each class. We can observe distinct clusters of symptoms, such as fever and cough; dyspnea, respiratory discomfort, and low $O_{2}$ saturation; loss of taste and loss of smell, that are more likely to occur in the same latent symptom profiles, which seem to correspond to symptoms with shared mechanisms. The estimated weights $\bm{\lambda}$ illustrate the changes in abundance of these latent classes over age, sex, and time. For example, among the deaths related to COVID-19, the weights of the fifth latent class are much larger in male deaths under 60 years old, compared to other demographic groups, whereas the sixth latent class has larger weights among deaths above 60 years old. Both symptom profiles include high probabilities of dyspnea, respiratory discomfort, and low $O_{2}$ saturation. The main difference in these two latent symptom profiles is that the former has a higher probability of observing fever and cough, whereas the latter has a higher probability of having at least one risk factor. These patterns are consistent with the explorative analysis in Figure 4.
@@
Figure 12: Posterior mean of latent parameters estimated in one synthetic dataset resampled from the COVID-19 surveillance data from Brazil. Panel (a): symptom profiles $\bm{\phi}$ for deaths related to COVID-19, ordered by the expected number of symptoms. Each column corresponds to one latent class, and the color correspond to the probability of observing the symptom within the corresponding latent class. Panel (b): latent class weights $\bm{\lambda}$ over sex, age, and time for deaths related to COVID-19. Panel (c) and (d) are the same parameters for deaths not related to COVID-19.